# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ II.2 Preliminaries II
Table 2: The distinct behaviors of various physical quantities across the Mott transition [50, 51, 54] from a Landau Fermi liquid (FL) to a Mott insulator (MI), and the transition [55, 56, 71] from a Landau FL to a composite Fermi liquid (CFL). In this work, we emphasize the different behaviors of the localization tensor (i.e., polarization fluctuations) and the leading-order scaling of charge fluctuations.
@@
where $\mathsf{G}^{ij}$ represents the real part of the localization tensor (i.e., polarization fluctuations) [59, 60, 61, 62]. Recently, $\mathsf{G}^{ij}$ has also been referred to as the “quantum weight” in Refs. [72, 73, 74]. In this work, we are interested in phase transitions from a Landau FL, where $\mathsf{G}^{ij}$ is divergent, to another phase where $\mathsf{G}^{ij}$ becomes finite. According to the general criterion in Ref. [59, 60, 61, 62], systems with a finite value of $\mathsf{G}^{ij}$ are classified as charge insulators.
@@ II.3 Summary of Results
If the above-mentioned phase transitions are indeed continuous, the critical points are expected to be incompressible, with a vanishing Drude weight and intriguing universal behaviors in the localization tensor and charge fluctuations, which will be summarized below. The descriptions of these continuous phase transitions are unified within a theoretical framework that introduces vortex excitations in Landau FLs, as will be discussed in Sec. III. Different charge insulators can then be realized by placing vortices in different states.
@@
Figure 3: (1) and (2) depict the predicted universal longitudinal resistivity jump $\Delta\rho_{xx}$ and Hall resistivity jump $\Delta\rho_{xy}$ at the transition between a Fermi liquid (FL) and a composite Fermi liquid (CFL) [56]; (3) illustrates the predicted universal longitudinal resistivity jump $\Delta\rho_{xx}$ at the transition between a Fermi liquid (FL) and a Mott insulator (MI) [51, 54]. In all of these phase transitions, the tuning parameter $g$ is related to the electron bandwidth, and the critical resistivity jumps are linked to the universal corner contribution of bipartite charge fluctuations, as described by Eq. (15).
@@
    Despite being compressible, CFLs follow the leading-order boundary-law scaling $\mathcal{F}\sim L$ of bipartite charge fluctuations, in contrast to the scaling $\mathcal{F}\sim L\log(L)$ seen in other compressible states like Landau FLs and superfluids. This difference arises because the instantaneous charge correlation from the gapless modes decays as $|\bm{x}|^{-5}$ (up to logarithmic corrections). Moreover, the gapped modes contribute a subleading corner term reminiscent of the behavior observed in gapped QH insulators [21]. Considering the subsystem geometry depicted in Fig. 1, the final result is
    =#$\mathcal{F}=\#L-\frac{\mathsf{G}}{\pi}\textrm{f}(\theta)+\ldots$ (12)
    where ##\## represents a non-universal constant, and $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ is the localization tensor introduced in Eq. (9) and Eq. (10). Based on the Halperin-Lee-Read (HLR) theory [63], we find the coefficient
@@
    =#$\mathcal{F}=\#L-\frac{C_{\rho}}{2}\textrm{f}(\theta)\log(L)+\ldots.$ (16)
@@
    The continuous Mott transition [51] can be detected by the distinct leading-order scalings of bipartite charge fluctuations: $\mathcal{F}^{c}\sim L\log(L)$ in the metallic phase and $\mathcal{F}^{c}\sim L$ in the insulating phase. At the non-conformal critical point, the localization tensor exhibits a universal divergence, as described in Eq. (14). The scaling of bipartite charge fluctuations $\mathcal{F}^{c}$ follows Eq. (16) with the coefficient $C_{\rho}$ given by the current central charge $C_{J}^{(\textrm{XY})}$ of the 3D XY universality class. The critical transport behavior is simpler due to the vanishing Hall response, and the relation in Eq. (15) to the critical resistivity jump still holds, with $\Delta\rho_{xy}=0$. The system also possesses a spin $\textrm{U}(1)$ symmetry, allowing for the definition of bipartite spin fluctuations $\mathcal{F}^{s}$. In both phases, the Fermi-surface states of spin excitations manifest as $\mathcal{F}^{s}\sim L\log(L)$. Richer physics can be revealed by generalizing the vortex theory in Sec. III to involve both charge and spin vortices. Further insights into the multicritical behaviors will be elaborated upon in Sec. V.2.
@@
    In the modified proposal [54] of the continuous Mott transition involving charge fractionalization, bipartite charge fluctuations $\mathcal{F}^{c}$ still exhibit distinct leading-order scalings of $\mathcal{F}^{c}\sim L\log(L)$ and $\mathcal{F}^{c}\sim L$ in the two phases. A key difference from the original theory [51] is observed in the universal behavior of $\mathcal{F}^{c}$ at the critical point, where it follows Eq. (16) but with a suppressed universal coefficient $C_{\rho}=2C_{J}^{(\textrm{XY})}/N^{2}$, where $C_{J}^{(\textrm{XY})}$ still represents the current central charge of the 3D XY fixed point. Here, $N$ is an even integer corresponding to the fractional electric charge $e/N$ carried by each charge carrier. Notably, in the time-reversal invariant version of Eq. (15) with $\Delta\rho_{xy}=0$, the critical resistivity jump $\Delta\rho_{xx}$ is enhanced by a factor of $N^{2}/2$, consistent with the experimentally observed large critical resistivity [52]. Another distinction from the original theory is seen in the bipartite spin fluctuations $\mathcal{F}^{s}$, which show boundary-law scaling $\mathcal{F}^{s}\sim L$ in the Mott insulator phase.
@@ III Vortex Theory Framework
For illustrative purposes, let us focus on the case of spinless electrons. The generalization to spinful cases follows in a similar way. Due to the $\textrm{U}(1)$ gauge redundancy in Eq. (17), both $b$ and $f$ are coupled to a dynamical $\textrm{U}(1)$ gauge field $a_{\mu}$, carrying equal and opposite charges of ±$\pm 1$. When a background field $A_{\mu}$ is introduced for the global $\textrm{U}(1)$ symmetry, the effective Lagrangian can be schematically expressed as
@@
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a]+\frac{\mathtt{i}}{2\pi}\tilde{a}\wedge\textrm{d}(A-a)+\mathcal{L}_{\textrm{vrtx}}[v,\tilde{a}]+\ldots$ (19)
@@
Central to our discussion is the electromagnetic response of the system. By definition, the conserved U(1) current in the vortex theory Eq. (19) is given by
@@
Therefore, the response function $\Pi^{\mu\nu}=\langle J^{\mu}J^{\nu}\rangle$ is determined by the fully dressed gauge-field propagator $\langle\tilde{a}_{\mu}\tilde{a}_{\nu}\rangle$. For the problem of a Fermi surface coupled to a gauge field $a_{\mu}=(a_{\tau},\bm{a})$, it is convenient to use the Coulomb gauge $\nabla\cdot\bm{a}=0$ [30, 31, 63, 78, 34]. At the level of the random phase approximation (RPA), Eq. (19) leads to the effective theory
$\mathcal{S}=\int_{k}\tilde{a}(-k)\frac{-\tilde{\Pi}(k)}{2}\tilde{a}(k)-\tilde{a}(-k)\Pi_{\textrm{CS}}(k)A(k)+\ldots$ (21)
where $k=(\omega,\bm{k})$ is a collective notion of frequency and momentum. The response kernels $\tilde{\Pi}$ and $\Pi_{\textrm{CS}}$ are $2\times 2$ matrices in the basis $\tilde{a}=(\tilde{a}_{\tau},\tilde{a}_{T})$ and $A=(A_{\tau},A_{T})$. In our convention (see App. B), the Chern-Simons kernel is $\Pi_{\textrm{CS}}(k)=\frac{|\bm{k}|}{2\pi}\sigma^{1}$, where $\sigma^{1}$ denotes the first Pauli matrix. The fully dressed $\tilde{\Pi}(k)$ contains contributions from both the fermion sector and the vortex sector
@@
where $\Pi_{f}$ is the response of the fermionic partons $f$ to $a_{\mu}$ and $\Pi_{v}$ is the response of the vortices $v$ to $\tilde{a}_{\mu}$. The electromagnetic response of electrons is then given by
@@
This is nothing but the Ioffe-Larkin rule Eq. (80) for the parton construction Eq. (17), where the duality relation in the chargon sector imposes $\Pi_{v}=-\Pi_{\textrm{CS}}\Pi_{b}^{-1}\Pi_{\textrm{CS}}$. To avoid redundant terminology, we refer to the equivalent expressions under duality transformations, Eq. (80), Eq. (23), and Eq. (22), as the Ioffe-Larkin rule [79].
From a different perspective, for any fermionic systems with a global U(1) symmetry, one can always introduce a gauge field $\tilde{a}_{\mu}$ to represent the conserved current $J^{\mu}$ via Eq. (20). Consequently, Eq. (21) can be interpreted as the bosonization [80, 81] of non-relativistic electrons.
@@
$\mathcal{W}_{\mathcal{C}}(\vartheta)=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}\tilde{a}\right),$ (24)
@@
electron $c$ spinless $c(\bm{x})=b(\bm{x})f(\bm{x})$ spinful $c_{\sigma}(\bm{x})=b(\bm{x})f_{\sigma}(\bm{x})$ spinful $c_{\sigma}(\bm{x})=b_{\sigma}(\bm{x})f_{\sigma}(\bm{x})$ldots
fermion $f$ Fermi surface Fermi surface Fermi surface Fermi surface
Table 3: The parton constructions $c\sim bf$ for the quantum phase transitions out of spinless and spin-$1/2$ Landau Fermi liquids at half-filling, where $\sigma=\uparrow,\downarrow$ represents the spin quantum number. The dual vortex $v$ is introduced in Eq. (19).
@@ IV Continuous CFL-FL Transition
We begin by discussing the critical Fermi surfaces for spinless electrons, focusing on the continuous transition between a Landau FL and a CFL at half-filling. The universal field theory of this transition was first explored in Ref. [55] and later refined in Ref. [56], with particular attention to its potential realization in twisted $\textrm{MoTe}_{2}$ bilayers through the tuning of the electron bandwidth. More recently, Ref. [71] reported that a CFL-FL transition is indeed realized in a microscopic model of spatially modulated Landau levels by adjusting the interaction strength, as evidenced by the vanishing and nonvanishing values of the Drude weight, although whether the transition is continuous remains to be confirmed.
In Sec. IV.1, we briefly discuss the vortex theory that connects Landau FL theory with HLR theory [63] and its relation to previous constructions [55, 56] through fermionic particle-vortex duality. In Sec. IV.2, we clarify the prediction of a universal constant term in Eq. (12) within the CFL phase, as derived from HLR theory, and the universal logarithmic term in Eq. (16) at the critical point, further examining its connection to the critical resistivity jump described in Eq. (15).

### v3

@@ II.2 Preliminaries II
Table 2: Distinct behaviors of various physical quantities across the transition [51, 52, 55] from a Landau Fermi liquid (FL) to a Mott insulator (MI), and the transition [56, 57, 81] from a Landau FL to a composite Fermi liquid (CFL). This work highlights the different behaviors of the localization tensor (i.e., polarization fluctuations) and the leading-order scaling of charge fluctuations.
@@
where $\mathsf{G}^{ij}$ represents the real part of the localization tensor (i.e., polarization fluctuations) [63, 64, 65, 66]. Recently, $\mathsf{G}^{ij}$ has also been referred to as the “quantum weight” in Refs. [82, 83, 84]. In this work, we are interested in phase transitions from a Landau FL, where $\mathsf{G}^{ij}$ is divergent, to another phase where $\mathsf{G}^{ij}$ becomes finite. According to the general criterion in Ref. [63, 64, 65, 66, 85], systems with a finite value of $\mathsf{G}^{ij}$ are classified as charge insulators.
@@ II.3 Summary of Results
If the above-mentioned phase transitions are indeed continuous, the critical points are expected to be incompressible, with a vanishing Drude weight and intriguing universal behaviors in the localization tensor and charge fluctuations, which will be summarized below. The descriptions of these continuous phase transitions are unified within a theoretical framework that introduces vortex excitations in Landau FLs, as will be discussed in Sec. III. Different charge insulators can then be realized by placing vortices in different quantum states.
@@
Figure 3: Panels (1) and (2) show the predicted universal jumps in longitudinal resistivity $\Delta\rho_{xx}$ and Hall resistivity $\Delta\rho_{xy}$ at the transition from a Fermi liquid (FL) to a composite Fermi liquid (CFL) [57]; panel (3) illustrates the predicted universal longitudinal resistivity jump $\Delta\rho_{xx}$ at the transition from a Fermi liquid (FL) to a Mott insulator (MI) [52, 55]. In all cases, the tuning parameter $g$ corresponds to the electron bandwidth. These resistivity jumps are directly connected to the universal corner contribution to bipartite charge fluctuations, as given in Eq. (15).
@@
    Despite being compressible, CFLs exhibit the leading-order boundary-law scaling $\mathcal{F}\sim L$ for bipartite charge fluctuations, in contrast to the $\mathcal{F}\sim L\log(L)$ scaling seen in other compressible phases like Landau FLs and superfluids. This difference arises because the instantaneous charge correlation from the gapless modes decays as $|\bm{x}|^{-5}$ (up to logarithmic corrections). In addition, the gapped modes give rise to a subleading corner term analogous to those found in gapped QH insulators [22]. For the subsystem geometry depicted in Fig. 1, the final result takes the form
    $\mathcal{F}=\#L-\frac{\mathsf{G}}{\pi}\textrm{f}(\theta)+\ldots$ (12)
    where #\# represents a non-universal constant, and $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ is the localization tensor introduced in Eq. (9) and Eq. (10). Based on the Halperin-Lee-Read (HLR) theory [67], we find the coefficient
@@
    $\mathcal{F}=\#L-\frac{C_{\rho}}{2}\textrm{f}(\theta)\log(L)+\ldots.$ (16)
@@
    The continuous Mott transition [52] can be detected through the distinct leading-order scalings of bipartite charge fluctuations: $\mathcal{F}^{c}\sim L\log(L)$ in the metallic phase and $\mathcal{F}^{c}\sim L$ in the insulating phase. At the non-conformal critical point, the localization tensor exhibits a universal divergence, as described in Eq. (14). The scaling of bipartite charge fluctuations $\mathcal{F}^{c}$ follows Eq. (16) with the coefficient $C_{\rho}$ given by the current central charge $C_{J}^{(\textrm{XY})}$ of the 3D XY universality class. Since $\Delta\rho_{xy}=0$, the resistivity jump is entirely longitudinal, and Eq. (15) remains applicable. The system also possesses a spin $\textrm{U}(1)$ symmetry, allowing for the definition of bipartite spin fluctuations $\mathcal{F}^{s}$. In both phases, the Fermi-surface states of spin excitations manifest as $\mathcal{F}^{s}\sim L\log(L)$. Richer physics can be revealed by extending the vortex theory in Sec. III to include both charge and spin vortices. Further insights into the associated multicritical behavior will be discussed in Sec. V.2.
@@
    In the modified proposal [55] of the continuous Mott transition involving charge fractionalization, the leading-order scaling of bipartite charge fluctuations $\mathcal{F}^{c}$ remains the same: $\mathcal{F}^{c}\sim L\log L$ in the metallic phase and $\mathcal{F}^{c}\sim L$ in the insulating phase. However, at the quantum critical point, the universal corner contribution in Eq. (16) is suppressed relative to the original theory [52], with $C_{\rho}=2C_{J}^{(\textrm{XY})}/N^{2}$, where $N$ is an even integer corresponding to the fractional electric charge $e/N$ carried by each charge carrier. Consequently, in the time-reversal-invariant version of Eq. (15) with $\Delta\rho_{xy}=0$, the critical longitudinal resistivity jump $\Delta\rho_{xx}$ is enhanced by a factor of $N^{2}/2$, consistent with the large critical resistivity observed experimentally [53]. Another key difference from the original theory is that the bipartite spin fluctuations $\mathcal{F}^{s}$ obey boundary-law scaling $\mathcal{F}^{s}\sim L$ in the Mott insulating phase.
@@ III Vortex Theory Framework
For illustrative purposes, let us focus on the case of spinless electrons. The generalization to spinful cases follows in a similar way. Due to the $\textrm{U}(1)$ gauge redundancy in Eq. (17), both $b$ and $f$ are coupled to a dynamical $\textrm{U}(1)$ gauge field $a_{\mu}$, carrying equal and opposite charges of $\pm 1$. When a background field $A_{\mu}$ is introduced for the global $\textrm{U}(1)$ symmetry, the effective Lagrangian can be schematically expressed as
@@
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a]+\frac{\mathtt{i}}{2\pi}\tilde{a}\wedge\textrm{d}(A-a)+\mathcal{L}_{\textrm{vrtx}}[v,\tilde{a}$$\ldots$ (19)
@@
electron $c$ spinless $c(\bm{x})=b(\bm{x})f(\bm{x})$ spinful $c_{\sigma}(\bm{x})=b(\bm{x})f_{\sigma}(\bm{x})$ spinful $c_{\sigma}(\bm{x})=b_{\sigma}(\bm{x})f_{\sigma}(\bm{x})$ $\ldots$
fermion $f$ Fermi surface Fermi surface Fermi surface Fermi surface
Table 3: Summary of the parton constructions for continuous quantum phase transitions out of spinless and spin-$1/2$ Landau Fermi liquids at half-filling. Here, $\sigma=\uparrow,\downarrow$ denotes the spin quantum number. The dual vortex field $v$ is introduced in Eq. (19).
III.1 Electromagnetic Response
The electromagnetic response of the system lies at the heart of our discussion. In the vortex theory Eq. (19), the conserved $\mathrm{U}(1)$ current is given by
@@ III.1 Electromagnetic Response
Consequently, the response function $\Pi^{\mu\nu}$ associated with $J^{\mu}$ (see App. B for its definition) is determined by the fully dressed gauge-field propagator $\langle\tilde{a}_{\mu}\tilde{a}_{\nu}\rangle$.
For the problem of a Fermi surface coupled to a gauge field $a_{\mu}=(a_{\tau},\bm{a})$, it is convenient to work in the Coulomb gauge $\nabla\cdot\bm{a}=0$ [31, 32, 67, 88, 35]. At the level of the random phase approximation (RPA), Eq. (19) leads to the effective theory (see App. B for our convention)
$\mathcal{S}=\int_{k}\tilde{a}(-k)\frac{\tilde{\Pi}(k)}{2}\tilde{a}(k)-\tilde{a}(-k)\Pi_{\textrm{CS}}(k)A(k)+\ldots$ (21)
where $k=(\omega,\bm{k})$ collectively denotes frequency and momentum. The response kernels $\tilde{\Pi}$ and $\Pi_{\textrm{CS}}$ are $2\times 2$ matrices expressed in the basis $\tilde{a}=(\tilde{a}_{\tau},\tilde{a}_{T})$ and $A=(A_{\tau},A_{T})$. In our convention, the Chern-Simons kernel takes the form $\Pi_{\textrm{CS}}(k)=-\sigma^{1}|\bm{k}|/(2\pi)$, where $\sigma^{1}$ denotes the first Pauli matrix. The full gauge kernel $\tilde{\Pi}(k)$ receives contributions from both the fermionic and vortex sectors
@@
where $\Pi_{f}$ is the response of the fermionic partons $f$ to $a_{\mu}$ and $\Pi_{v}$ is the response of the vortices $v$ to $\tilde{a}_{\mu}$. The resulting gauge-invariant electromagnetic response of the electron system is then
@@
This general relation holds across the entire phase diagram, provided gauge fluctuations are not too strong.
This expression is equivalent to the Ioffe-Larkin rule Eq. (88) for the original parton construction Eq. (17), where the duality relation in the chargon sector imposes $\Pi_{v}=-\Pi_{\textrm{CS}}\Pi_{b}^{-1}\Pi_{\textrm{CS}}$. To avoid redundant terminology, we refer to the equivalent expressions under duality transformations, Eq. (88), Eq. (23), and Eq. (22), as the Ioffe-Larkin composition rule [89].
From a different perspective, for any fermionic systems with a global U(1) symmetry, one can always introduce a gauge field $\tilde{a}_{\mu}$ to represent the conserved current $J^{\mu}$ via Eq. (20). Consequently, Eq. (21) can be interpreted as the “bosonization” of non-relativistic electrons [90, 91].
@@
[…]
@@
III.2 Quantum Criticality
In this subsection, we clarify the underlying minimal conditions necessary for the universal behaviors of the localization tensor Eq. (14) and the bipartite fluctuations Eq. (16) to manifest at critical Fermi surfaces.
As will become evident from the examples in the following sections, both Eq. (14) and Eq. (16) follow directly from the form of the critical response function
$\Pi^{-1}=\Pi_{\textrm{FS}}^{-1}+\Pi_{\textrm{CFT}}^{-1},$ (26)
where $\Pi_{\textrm{FS}}$ and $\Pi_{\textrm{CFT}}$ denote the response functions of a Fermi surface and a CFT, respectively. This structure appears as a natural consequence of the Ioffe-Larkin rule, Eq. (88) or Eq. (23). However, it also relies on an additional requirement: that the critical behavior of the chargon (or vortex) sector remains unaffected by its coupling to other low-energy degrees of freedom—a point that requires justification.
At the critical point, the effective theory Eq. (18) can be schematically written as
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a]+\mathcal{L}_{\textrm{CFT}}[b,A-a]+\mathcal{L}[a]+\mathcal{L}_{\textrm{cp}}[O_{f},O_{b}]+\ldots$ (27)
The two parton sectors are coupled both directly via the term $\mathcal{L}_{\textrm{cp}}$, which involves gauge-invariant operators $O_{f}$ and $O_{b}$, and indirectly through their mutual coupling to the dynamical gauge field $a_{\mu}$. A key task is to determine the fate of these couplings at the quantum critical point.
The leading direct coupling is expected to have the simple form $\mathcal{L}_{\textrm{cp}}\propto O_{f}O_{b}$, with $O_{f}=f^{\dagger}f$ and $O_{b}=|b|^{2}$. It is well known [92] that coupling a bosonic mode to a Fermi surface induces a Landau damping term of the form $\propto\int_{\omega,\bm{k}}\frac{|\omega|}{|\bm{k}|}|O_{b}(\omega,\bm{k})|^{2}$, which is irrelevant if the scaling dimension of $O_{b}$ in the CFT satisfies $\Delta[O_{b}]>3/2$. This condition is firmly established for the 3D XY transition, as confirmed by experiments, Monte Carlo simulations, and conformal bootstrap (see Ref. [93]). For the continuous transition between a bosonic Laughlin state at $\nu=1/2$ and a superfluid, this condition is also suggested by large-$N$ analysis [86, 56].
To understand the IR behavior of the dynamical gauge field $a_{\mu}$, it is essential to incorporate the effects of Landau damping, which arises from its coupling to the Fermi surface [51, 52]. The effective action for the transverse component of the gauge field takes the form
$\mathcal{S}_{\textrm{eff}}[a]=\int_{\omega,\bm{k}}\left(\gamma\frac{|\omega|}{|\bm{k}|}+\chi|\bm{k}|^{2}+\ldots\right)|a(\omega,\bm{k})|^{2},$ (28)
where $\chi$ denotes the diamagnetic susceptibility of the $f$-fermions, and the ellipsis includes other contributions such as the polarizability of the $b$-bosons. From this expression, it is clear that gauge fluctuations do not affect the critical properties of the chargon (or vortex) sector. Due to the identical scaling of $\omega$ and $\bm{k}$ in the CFT, the Landau damping term acts effectively as a Higgs mass term and quenches the dynamics of the gauge field. This suppression of gauge fluctuations lends further support to the validity of the Ioffe-Larkin rule.
In summary, we have reviewed the justification for dynamical decoupling [51, 52, 56, 57, 55]: both the direct couplings between the two parton sectors and the gauge couplings in $\mathcal{L}_{\textrm{CFT}}[b,A-a]$ are irrelevant at the critical point [^4] These couplings, however, become important for describing crossovers away from the critical point and for understanding the full phase diagram. For a detailed introduction to the phenomenology, see Ref. [51, 52].. This leads to the important conclusion that correlation functions in the chargon (or vortex) sector retain their CFT form.
We therefore argue that the universal behaviors in Eq. (14) and Eq. (16) can be understood as direct consequences of two essential ingredients: (1) the Ioffe-Larkin rule and (2) dynamical decoupling. These conditions are expected to be satisfied in the examples listed in Tab. 3, which will be discussed in detail in Sec. IV and Sec. V, and are likely to apply to other similar continuous transitions out of Landau FLs as well.
@@ IV Continuous CFL-FL Transition
We begin by discussing the critical Fermi surfaces for spinless electrons, focusing on the continuous transition between a Landau FL and a CFL at half-filling. The universal field theory for this transition was first proposed in Ref. [56] and subsequently refined in Ref. [57], with particular emphasis on its potential realization in moiré materials through tuning the electron bandwidth. Experimental evidence supporting the existence of zero-field CFL phases, as well as signatures of the CFL-FL transition, has recently been observed in twisted $\textrm{MoTe}_{2}$ bilayers [60, 61] and in multilayer graphene systems [62]. In parallel, Ref. [81] reported the realization of a CFL-FL transition in a microscopic model of spatially modulated Landau levels by varying the interaction strength, as indicated by the vanishing and nonvanishing behavior of the Drude weight, although whether the transition is truly continuous remains to be determined.
In Sec. IV.1, we briefly discuss the vortex theory that connects Landau FL theory with the HLR theory [67], as well as its relation to the earlier construction [56, 57] through fermionic particle-vortex duality. In Sec. IV.2, we clarify the prediction of a universal constant term in Eq. (12) within the CFL phase, as derived from HLR theory, and the universal logarithmic term in Eq. (16) at the critical point, and we further examine their connection to the critical resistivity jump described in Eq. (15).